# When Should Selfish Miners Double-Spend?

### v1

@@ 4 Revenue & Double-Spending of $L$-Stubborn Mining
To find the optimal $L$, one could evaluate $\rho_{L}$ in (10) and find the $L$ that provides the highest revenue. As the calculations involve binomial coefficients and (10) involves fractions, it is not straightforward to find $L^{*}$. Instead, we provide a much simpler method to obtain the optimal $L$. To do so, we resort to the linearization method of the MDP problem [6] and use it in a narrow-sense. Note that, every time a block is mined, the adversary faces a decision as to whether it should publish $A$-chain and end the cycle, or continue. Clearly, if $A-H>1$, the decision should be to continue the attack cycle as the adversary can gain additional advantage at no-cost in these situations. On the other hand, if $A=H+1$, the risk is imminent as the adversary loses the advantage of overwriting the $H$-chain. As a result, we can narrow down the decision problem to a specific case where $A=H+1$ and analyze the situation therein via the linearization method.
To that end, we define
$w_{\rho}(x,y)=(1-\rho)x-\rho y,$ (19)
as in [6], and $u$ and $v_{a}$ as
@@
Next, using the definitions above and intermediary results in Appendix A, we provide the following quasiconcavity result related to the $L$-stubborn mining.
@@
Theorem 3 essentially says that $\rho_{L}$ is quasiconcave for integer $L$, which implies if $L^{*}>k+1$, $(k+1)$-stubborn mining is still more profitable than honest mining. We relegate the proof of Theorem 3 to the Appendix A as it requires some intermediary definitions and results. These results also shed light on a simple algorithm that can be used to find $L^{*}$.
Next, we provide Algorithm 1 that finds the optimal $L^{*}$ of $L$-stubborn mining. For that, let $\rho_{1}$, $\rho_{2}$ and $\rho_{\infty}$ be defined as follows
$\rho_{1}$ $=\alpha,$ (21)
$\rho_{2}$ $=\frac{\alpha\beta^{2}(4\alpha+\gamma(1-2\alpha))-\alpha^{3}}{1-\alpha(1+(2-\alpha)\alpha)},$ (22)
$\rho_{\infty}$ $=\frac{\alpha}{\beta}-\frac{(1-2\alpha)(1-\gamma)}{\beta\gamma}(1-\beta C((1-\gamma)\alpha\beta)),$ (23)
which are the revenues of honest mining, selfish mining [4], and equal-fork stubborn mining [7, 28], respectively. Note that each of these values can be obtained by (10), where for $\rho_{\infty}$, $L_{s}$ can be set to zero as the $H$-chain will catch the $A$-chain eventually, i.e., no attack ends successfully, but revenue comes from matching. In Appendix A, we use the intermediary results to show that Algorithm 1 below outputs $L^{*}$.
@@
4.2 Combined Revenue
In this section, we provide a formula that incorporates the profit of double-spending into the revenue the adversary makes from $(k+1)$-stubborn mining. To do so, we make the assumption of [9], i.e., we assume that the profit of every successful double-spend is equal to $\mathcal{R}$ units of block rewards. Further, as in [9], we consider the replacement of every confirmed block as different instances of double-spend events, i.e., a cycle where $l$ honest blocks are confirmed and later replaced, result in extra $l\cdot\mathcal{R}$ revenues in block reward units. Notice that, in this paper, we treat the revenue as the relative adversarial block rewards in the system. Thus, we also treat the double-spending revenue as extra block rewards the adversary can get in addition to the actual block rewards. Other considerations are left as simple extensions which the reader can easily calculate given the approaches we provide here.
@@ 4.2 Combined Revenue
(Combined Revenue) The combined revenue of $(k+1)$-stubborn mining and the associated double-spending is given as
@@
The reasoning for this is as follows: Out of every $\left((k+1)_{s}+(k+1)_{u}\right)$ blocks that make it into the prefix chain, if the adversary followed the honest protocol, $\alpha$ proportion would be its profit. However, with the $(k+1)$-stubborn mining protocol, $(k+1)_{s}+(k+1)_{u,a}$ blocks make it into the prefix chain. Further, the double-spending results in $V$ extra rewards for the adversary. However, if only Service happens, $F-V$ amount of rewards is lost.
5 $S$-Stealth Mining
[…]
In $S$-stealth mining, the strategy of keeping track of the length of $A$-chain is still the same, however, matching only happens if $A=H=S-1$. In other words, whenever the honest miners mine a new block on a height $H<S-1$, the adversary does not match the honest block. On the other hand, if $A=S-1$ and $H=S-2$ and the next block is honest on height $S-1$, the adversary matches this block.
@@ 5 $S$-Stealth Mining
(Revenue) The revenue of $S$-stealth mining is given as
@@
Proof:  For a successful attack, we follow the same arguments provided in the proof of Theorem 1. When an attack is unsuccessful, adversarial blocks make it into the offset chain only if the $H$-chain has reached length $S-1$, and the next honest block on height $S$ switches its prefix (w.p. $\gamma$) to the $A$-chain which has length $S-1$. $\blacksquare$
@@
Theorem 6
[…]
Notice that $(k+1)$-stealth mining is essentially the private attack of Nakamoto [1] where the adversary gives up pursuing the attack if it falls behind before reaching length $k$. If it reaches length $k$ but the honest chain catches it at length $k$, the network split $\gamma$ could still help the adversary to double-spend. However, as matching only happens at length $k$, there is no Move-Funds event possible in $(k+1)$-stealth mining.
The formula we provide in Theorem 5 can be used to determine the revenue of $(k+1)$-stealth mining which is no different than the private attack with no-trail. If $\sigma_{k+1}\geq\alpha$, this would imply that trying private attack with no-trail comes at no-cost, i.e., mining revenues are still above or equal to $\alpha$. Hence, similar to the case of $L$-stubborn mining we define
$S^{*}$ $=\arg\max_{S}\sigma_{S},$ (35)
$\bar{S}$ $=\sup\{S:\sigma_{S}\geq\alpha\},$ (36)
and provide their associated values for given $\alpha$ and $\gamma$ in the numerical results section. For the sake of completeness, we provide here the quasiconcavity of $S$-stealth mining and an algorithm to find $S^{*}$. The proofs are provided in Appendix B.
@@
5.1 Combined Revenue
Here, we first provide the approach where each honest block that is confirmed but does not make it into the offset chain, results in $\mathcal{R}$ additional block rewards for the adversary. In this case, the revenue is provided in Theorem 8.
@@ 5.1 Combined Revenue
(Combined Revenue) The combined revenue of $(k+1)$-stealth mining and the associated double-spending is given as
@@ 6 Numerical Results
We start by evaluating $L^{*}$ for each value of $\alpha$ and $\gamma$ using Algorithm 1. The results are provided in Figure 2(a). Note that the boundary between $L^{*}=1$ and $L^{*}=2$, i.e., the honest mining and selfish mining, was already provided in [4]. Our results here show new boundaries by providing regions where waiting for the private branch to extend further is beneficial in terms of the revenue. Next, in Figure 2(c) for each $\alpha$ and $\gamma$, we provide the relative revenue of optimal $L^{*}$-stubborn mining attack compared to honest mining. For all values of $\gamma$ the relative revenue starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$. However, as $\gamma$ increases, the revenue increases faster initially. Finally, in Figure 2(e), we provide $\bar{L}$ for each $\alpha$ and $\gamma$, which we believe has a significant meaning related to $(k+1)$-confirmation rule. As we proved that $\rho_{L}$ has a quasiconcave behavior, if $\bar{L}>k$, then $(k+1)$-stubborn mining is more profitable than honest mining, hence following the $(k+1)$-stubborn mining allows double-spending at no-cost. This in turn means, for each $\alpha$ and $\gamma$, $k$-confirmation rule has to be set according to $k\geq\bar{L}(\alpha,\gamma)$ to minimize the risk of double-spending that comes at no-cost to the attacker.
@@
Similar to the results of $L$-stubborn mining, Figure 2 also contains related results for $S$-stealth mining. Observing the $S^{*}$ provided in Figure 2(b), it is clear that $S^{*}\leq L^{*}$ which is straightforward, but also $S^{*}\leq 4$ for $\alpha<0.5$ and $\gamma<1$. In other words, in stealth mining, adversary cannot act in a stubborn manner forever, whereas stubborn mining is the best strategy for selfish mining in the white region of Figure 2(a). In Figure 2(d), we provide the relative revenue of optimal $S^{*}$-stealth mining attack compared to honest mining. Although it starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$ similar to the $L^{*}$-stubborn mining, the rise in revenue is not as steep initially. Further, as expected, the revenue difference between $L^{*}$-stubborn mining and $S^{*}$-stealth mining increases with $\gamma$, as matching is only present in the last step in stealth mining. Finally, Figure 2(f) shows the values of $\bar{S}$, where as $\sigma_{\infty}=0$, we have $\bar{S}<\infty$ for all values of $\alpha<0.5$ and $\gamma<1$.
[…]
@@
For $(k+1)$-stubborn mining, the double-spending risk for $k=6$, $\alpha=0.41$ and $\gamma=0$ is $0.092$. For $k=6$ and $\gamma=0$, as $\alpha$ approaches $0.5$, the double-spending risk approaches $0.21$. On the other hand, for $k=6$, $\alpha=0.41$ and $\gamma=1$, the double-spending risk is $0.002$, whereas service probability is $0.59$. For $\gamma=0$, the same values apply to $(k+1)$-stealth mining as well. On the other hand, for $(k+1)$-stealth mining with $k=6$, $\alpha=0.41$ and $\gamma=1$, the double-spending risk is increased to $0.108$, whereas service probability increases to $0.892$. As mentioned earlier, the cost is a cut in relative revenue, which drops from $\frac{\rho_{7}}{\alpha}=1.639$ to $\frac{\sigma_{7}}{\alpha}=1.09$.
[…]
7 Conclusion and Extensions

### v2

@@ 4.1 Optimizing the Revenue Ratio
As we mentioned earlier, our $L$-stubborn mining strategy is a generalization and extension of the existing mining strategies of honest mining $(L=1)$, selfish mining [4] $(L=2)$, equal fork stubborn mining [6] $(L=\infty)$. As a result, for a given $(\alpha,\gamma)$ pair, a natural question is, what integer value of $L$ maximizes the revenue ratio $\rho_{L}$. To find the optimal $L$, one could evaluate $\rho_{L}$ in (18) and find the $L$ that provides the highest revenue ratio. As the calculations involve binomial coefficients and (18) involves fractions, it is not straightforward to find $L^{*}$. Instead, here, we provide a much simpler method to obtain the optimal $L$. To do so, we resort to the linearization method of the MDP problem [5] which is presented in Appendix A, and use it in a narrow-sense. Note that, every time a block is mined, the adversary faces a decision as to whether it should publish $A$-chain and end the cycle, or continue. Clearly, if AH>1A-H>1, the decision should be to continue the attack cycle as the adversary can gain additional advantage at no-cost in these situations. On the other hand, if $A=H+1$, the risk is imminent as the adversary loses the advantage of overriding the $H$-chain if the next arrival is honest. As a result, we can narrow down the decision problem to a specific case where $A=H+1$ and analyze the situation therein via the linearization method.
To that end, we define $u$ and $v_{a}$ as
@@
Next, using the definitions above and intermediary results in Appendix B, we provide the following quasiconcavity result related to the $L$-stubborn mining.
@@
Theorem 2 essentially says that $\rho_{L}$ is quasiconcave for integer $L$. We relegate the proof of Theorem 2 to the Appendix B as it requires some intermediary definitions and results. These results also shed light on a simple algorithm that can be used to find $L^{*}$.
Next, we provide Algorithm 1 that finds the optimal $L^{*}$ of $L$-stubborn mining. For that, let $\rho_{1}$, $\rho_{2}$ and $\rho_{\infty}$ be defined as in Definition 5, which are the revenue ratios of honest mining, selfish mining [4], and equal-fork stubborn mining [6, 37], respectively. Note that each of these values can be obtained by (18), where for $\rho_{\infty}$, $L_{s}$ can be set to zero as the $H$-chain will catch the $A$-chain eventually, i.e., no attack ends successfully, but the revenue comes from matching. In Appendix B, we use the intermediary results to show that Algorithm 1 outputs $L^{*}$, which essentially solves the decision problem mentioned earlier.
@@
4.2 $(k+1)$-Stubborn Mining and Double-Spending
In the analysis of double-spending, we focus on $L=(k+1)$-stubborn mining strategy for the ease of the analysis. However, we note that, the choice of $L$ does not need to equal $k+1$ for the adversary to get free shots at double-spending. For example, the attacker can try to use the optimal value $L=L^{*}$ that maximizes the revenue ratio of coinbase rewards and still some attack cycles result in double-spending. More specifically, if
  1. 1.
    $L<k+1$, successful attack cycles that reach $A=k+1$ may still result in double-spending.
  2. 2.
    $L>k+1$, both successful and unsuccessful attack cycles may result in double-spending.
We do not focus on the double-spending analysis of these two cases to avoid a complicated analysis. We pick $L=k+1$ when focusing on the double-spending since under the $(k+1)$-stubborn strategy the aim of the attacker is to undo an $H$ chain with length $H\geq k$ whenever possible. As a result, $(k+1)$-stubborn mining can be seen as a sustained double-spending attack until it succeeds in double-spending. For such an attack, we already provided the formula to calculate the revenue ratio of the $(k+1)$-stubborn mining. Here, we provide the probability that an attack cycle results in Double-Spending event.
Note that, even when $L=k+1$, a successful attack cycle does not immediately imply a successful double-spending. It only implies that the $A$-chain will make it into the offset chain eventually. Similarly, an unsuccessful attack cycle does not immediately imply that $tx$ will be confirmed. Next, we provide the probability of each event regarding $tx$ and $tx^{\prime}$ for $(k+1)$-stubborn mining with $k$-confirmation rule.
@@ 4.2 $(k+1)$-Stubborn Mining and Double-Spending
(Double Spend) Under $k$-confirmation rule, each attack cycle of $(k+1)$-stubborn mining ends with Service, Move-Funds or Double-Spending with probabilities,
$P(\textbf{Double-Spending})$ $=\sum_{m=0}^{k}P_{s}(k+1,m)(1-\gamma)^{m-\mathbbm{1}_{m=k}}+P_{u}(k)(1-\gamma)^{k-1}\gamma,$ (23)
$P(\textbf{Service})$ $=\sum_{m=0}^{k}P_{u}(m)(1-\gamma)^{m},$ (24)
$P(\textbf{Move-Funds})$ $=1-P(\textbf{Double-Spending})-P(\textbf{Service}).$ (25)
Here, we remind the reader that $\rho_{k+1}$ is the revenue ratio of the pool when it employs the $(k+1)$-stubborn mining strategy and $\bar{L}=\sup\{L:\rho_{L}\geq\alpha\}$, i.e., the largest value of $L$ such that the revenue ratio of $L$-stubborn mining is higher than the honest mining strategy. Hence, due to the quasiconcavity proven in Theorem 2, under any $k$-confirmation rule with $k<\bar{L}$ is compromised by $(k+1)$-stubborn mining strategy since the adversary gains more revenue ratio by employing this strategy compared to the revenue ratio of the honest mining strategy, i.e., $\rho_{k+1}\geq\alpha$, and at the same time it can double-spend at no-cost. In other words, for a given $\alpha$ and $\gamma$, $k$-confirmation rule should be at least $k\geq\bar{L}$ to make sure that double-spending attack has loss in revenue ratio. Note that, if $\rho_{\infty}\geq\alpha$, no $k$-confirmation is safe as double-spending has zero-cost. From this perspective, the safety of a blockchain system fails well before adversarial proportion reaches $\alpha$. We investigate this phenomenon further in the numerical results section by providing $L^{*}$ and $\bar{L}$ for each $\alpha$ and $\gamma$.
4.3 Combined Revenue Ratio
In this section, we provide a formula that incorporates the profit of double-spending into the revenue ratio the adversary makes from $(k+1)$-stubborn mining. To do so, we make the assumption of [8], i.e., we assume that the profit of every successful double-spend is equal to $\mathcal{R}$ units of block rewards. Further, as in [8], we consider the replacement of every confirmed block as different instances of double-spend events, i.e., a cycle where $l$ honest blocks are confirmed and later replaced, result in extra $l\cdot\mathcal{R}$ revenues in block reward units. Thus, we also treat the double-spending revenue as extra block rewards the adversary can get in addition to the actual block rewards. Other considerations are left as simple extensions which the reader can easily calculate given the approaches we provide here.
Theorem 4
(Combined Revenue Ratio) The combined revenue ratio of $(k+1)$-stubborn mining and the associated double-spending is given as
@@ 4.3 Combined Revenue Ratio
The reasoning for this is as follows: Out of every $\left((k+1)_{s}+(k+1)_{u}\right)$ blocks that make it into the prefix chain, if the adversary followed the honest protocol, $\alpha$ fraction would be its revenue ratio. However, with the $(k+1)$-stubborn mining protocol, $(k+1)_{s}+(k+1)_{u,a}$ blocks make it into the prefix chain. Further, the double-spending results in $V$ extra rewards for the adversary. However, if only Service happens, FVF-V amount of rewards is lost.
5 Analysis of $S$-Stealth Mining
An interesting observation of the results in Theorem 3 is the fact that $P(\textbf{Double-Spending})$ is decreasing in $\gamma$ due to the fact that, even if the attack cycles end successfully, $\gamma$ increases the probability that the prefix of $H$-chain switches to $tx^{\prime}$ before $tx$ is confirmed, because the adversary matches each honest block before $A=L$. Thus, for a given $\alpha$, while $\gamma$ increases, both $P(\textbf{Double-Spending})$ and $P(\textbf{Service})$ decrease and $P(\textbf{Move-Funds})$ increases. Further, even if the $H$-chain still contains $tx$ in its prefix, every time an honest block is mined, the adversary matches it, i.e., releases a competing block on the same height which contains $tx^{\prime}$ in its prefix, which would make the merchant providing services $tx$ suspicious. In this section, we provide a modified version of $L$-stubborn mining, which we call $S$-stealth mining, aiming to make the detection of the double-spending attack difficult and to increase $P(\textbf{Double-Spending})$ in each attack cycle.
In $S$-stealth mining, the strategy of keeping track of the length of $A$-chain is still the same, however, matching only happens if A=H=S1A=H=S-1. In other words, whenever the honest miners mine a new block on a height H<S1H<S-1, the adversary does not match the honest block. On the other hand, if A=S1A=S-1 and H=S2H=S-2 and the next block is honest on height S1S-1, the adversary matches this block.
@@ 5 Analysis of $S$-Stealth Mining
(Revenue Ratio) The revenue ratio of $S$-stealth mining is given as
@@
Proof:  For a successful attack, we follow the same arguments provided in the proof of Theorem 1. When an attack is unsuccessful, adversarial blocks make it into the offset chain only if the $H$-chain has reached length S1S-1, and the next honest block on height $S$ switches its prefix (w.p. $\gamma$) to the $A$-chain which has length S1S-1.  $\blacksquare$
@@
5.1 Optimizing the Revenue Ratio
Here, we first provide the theorem stating the quasiconcavity of $S$-stealth mining. Even though the goal in $S$-stealth mining is to increase the double-spending chance instead of focusing on the revenue ratio, for the sake of completeness, for a given $(\alpha,\gamma)$ pair, we also provide an algorithm to find $S^{*}$, i.e., the integer value of $S$ maximizing the revenue ratio. The proofs are provided in Appendix C.
@@ 5.1 Optimizing the Revenue Ratio
5.2 $(k+1)$-Stealth Mining and Double-Spending
Similar to the $(k+1)$-stubborn mining, $(k+1)$-stealth mining can also be seen as a sustained double-spending attack until it succeeds in double-spending. The difference is that the attacker here gives up some coinbase revenue by avoiding the matching actions, which allows the attack to be not exposed as well as increases the double-spending probability. Notice that $(k+1)$-stealth mining is essentially the private attack of Nakamoto [1] with the exception that, here, the adversary gives up pursuing the attack if it falls behind before reaching length $k$ to avoid revenue losses. If it reaches length $k$ but the honest chain catches it at length $k$, the network split $\gamma$ could still help the adversary to double-spend. However, as matching only happens at length $k$, there is no Move-Funds event possible in $(k+1)$-stealth mining. We already provided the formula to calculate the revenue ratio of the $k+1$-stealth mining. Here, we provide the probability that an attack cycle results in Double-Spending event.
@@ 5.2 $(k+1)$-Stealth Mining and Double-Spending
(Double Spend) Under the $k$-confirmation rule, each attack cycle of $(k+1)$-stealth mining ends with Service or Double-Spending with probabilities
$P(\textbf{Double-Spending})$ $=\sum_{m=0}^{k}P_{s}(k+1,m)+P_{u}(k)\gamma,$ (36)
$P(\textbf{Service})$ $=1-P(\textbf{Double-Spending}).$ (37)
The formula we provide in Theorem 5 can be used to determine the revenue ratio of $(k+1)$-stealth mining, i.e., $\sigma_{k+1}$, which is no different than the private attack with no-trail. If $\sigma_{k+1}\geq\alpha$, this would imply that trying private attack with no-trail comes at no-cost, i.e., the revenue ratio is still above or equal to $\alpha$. Hence, similar to the case of $L$-stubborn mining, we remind the reader that $\bar{S}=\sup\{S:\sigma_{S}\geq\alpha\}$, i.e., the largest value of $S$ such that the revenue ratio of $S$-stealth mining is higher than the honest mining strategy. Hence, due to the quasiconcavity proven in Theorem 6, under any $k$-confirmation rule with $k<\bar{S}$ is compromised by $(k+1)$-stealth mining strategy since the adversary gains more revenue ratio by employing this strategy compared to the revenue ratio of the honest mining strategy, i.e., $\sigma_{k+1}\geq\alpha$, and at the same time it can double-spend at no-cost. We provide the associated values of $S^{*}$ and $\bar{S}$ for given $\alpha$ and $\gamma$ in the numerical results section.
5.3 Combined Revenue Ratio
Here, we first provide the approach where each honest block that is confirmed but does not make it into the offset chain, results in $\mathcal{R}$ additional block rewards for the adversary. In this case, the revenue ratio is provided in Theorem 8. The metric introduced in Theorem 8 is similar to the subversion gain defined as in [10].
Theorem 8
(Combined Revenue) The combined revenue ratio of $(k+1)$-stealth mining and the associated double-spending is given as
@@ 5.3 Combined Revenue Ratio
Note that, $R^{*}_{k}$, defined here, is closely related to the subversion bounty defined in [10, Section IV.B.2]. As matched honest blocks are not considered for double-spending since they can make the merchant suspicious, subversion gain and subversion bounty of stealth mining is much closer to the definition of [10] than the corresponding results provided for stubborn mining in the previous section.
@@
$\alpha$ $\rho_{2}$[4] $\rho_{L^{*}}$ $\rho_{OSM}$[5] $\rho_{PTO}$[9]
1/3 0.33269 0.33333 0.33705 0.33705
0.35 0.36651 0.36651 0.37077 0.37077
0.375 0.42118 0.42118 0.42600 0.42600
0.4 0.48372 0.48372 0.48866 0.48866
0.425 0.55801 0.55801 0.56808 0.56809
0.45 0.65177 0.66248 0.66891 0.66894
0.475 0.78255 0.80043 0.80172 0.80184
Table 1: Comparison between the revenue ratio of $L$-Stubborn mining and MDP models [9, 5] for $\gamma=0$.
@@ 6 Numerical Results
We start by evaluating $L^{*}$ for each value of $\alpha$ and $\gamma$ using Algorithm 1. The results are provided in Figure 2(a). Note that the boundary between $L^{*}=1$ and $L^{*}=2$, i.e., the honest mining and selfish mining, was already provided in [4]. Our results here show new boundaries by providing regions where waiting for the private branch to extend further is beneficial in terms of the revenue ratio. Next, in Figure 2(c) for each $\alpha$ and $\gamma$, we provide the normalized revenue ratio of optimal $L^{*}$-stubborn mining attack compared to honest mining. For all values of $\gamma$ the normalized revenue ratio starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$, as expected since at $\alpha=0.5$, the adversary virtually creates all the blocks in the offset chain. However, as $\gamma$ increases, the ratio increases faster initially.
In Table 1, we compare the optimal revenue ratio obtained from $L^{*}$-stubborn mining with the revenue ratio obtained from selfish mining [4] (where $L=2$) and MDP models of Optimal Selfish Mining (OSM)[5] and Probabilistic Termination Optimization (PTO) [9] to demonstrate that our model performs close to the $\epsilon$-optimal selfish mining strategies. We use the data from [9, Table 1], where $\gamma=0$ and both MDP models are limited to maximum fork length $95$. It is easy to verify that our $L$-stubborn strategy behaves close to the optimal solution of MDP models for $\gamma=0$. The small difference can be attributed to the fact that, in MDP models, when the adversarial chain falls behind honest chain while having a length close to but smaller than $L$, adversary does not immediately adopt the honest chain but instead tries to catch up the honest chain which brings some additional revenue, albeit rather small.
Finally, in Figure 2(e), we provide $\bar{L}$ for each $\alpha$ and $\gamma$, which we believe has a significant meaning related to $(k+1)$-confirmation rule. As we proved that $\rho_{L}$ has a quasiconcave behavior, if $\bar{L}>k$, then $(k+1)$-stubborn mining is more profitable than honest mining, hence following the $(k+1)$-stubborn mining allows double-spending at no-cost. This in turn means, for each $\alpha$ and $\gamma$, $k$-confirmation rule has to be set according to $k\geq\bar{L}(\alpha,\gamma)$ to minimize the risk of double-spending that comes at no-cost to the attacker.
@@
Similar to the results of $L$-stubborn mining, Figure 2 also contains related results for $S$-stealth mining. Observing the $S^{*}$ provided in Figure 2(b), it is clear that $S^{*}\leq L^{*}$ which is straightforward from the analysis provided earlier, but also $S^{*}\leq 4$ for $\alpha<0.5$ and $\gamma<1$. In other words, in stealth mining, adversary cannot act in a stubborn manner forever, whereas acting stubborn no matter what the length of the adversarial branch is the best strategy to maximize the revenue ratio for $L$-stubborn mining in the white region of Figure 2(a). In Figure 2(d), we provide the normalized revenue ratio of optimal $S^{*}$-stealth mining attack. Although it starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$ similar to the $L^{*}$-stubborn mining, the rise in the revenue ratio is not as steep initially. Further, as expected, the revenue difference between $L^{*}$-stubborn mining and $S^{*}$-stealth mining increases with $\gamma$, as matching is only present in the last step in stealth mining. Finally, Figure 2(f) shows the values of $\bar{S}$, where we have $\bar{S}<\infty$ for all values of $\alpha<0.5$ and $\gamma<1$, since $\sigma_{\infty}=0$.
In Figure 3 we provide the probabilities of the events Double-Spending, Move-Funds and Service for each attack cycle of $(k+1)$-stubborn mining where we pick $k=6$ as in Bitcoin. We also provide normalized revenue ratio of $(k+1)$-stubborn mining in Figure 3(d) where the red curve highlights the boundary where the ratio rises above 1. It is clear that the red curve is the same as the boundary between $\bar{L}=6$ and $\bar{L}=7$ provided in Figure 2(e) by definition. For all values above and right of the red curve, a transaction is at risk of getting reversed in the sense that it is either not going to be confirmed at all or worse, it will be confirmed and then replaced with a conflicting transaction at no-cost to an adversary. The associated risks are provided in Figure 3(b) and Figure 3(a), respectively. As it is clear from Theorem 3, $P(\textbf{Double-Spending})$ is decreasing in $\gamma$ due to the fact that even if the attack cycles end successfully, $\gamma$ increases the probability of Move-Funds event as the prefix of the $H$-chain switches to $tx^{\prime}$ before $tx$ is confirmed because the adversary matches each honest block before $A=L$. Hence, for a given $\alpha$ as $\gamma$ increases, both $P(\textbf{Double-Spending})$ and $P(\textbf{Service})$ decrease, whereas $P(\textbf{Move-Funds})$ increases. Further, as mentioned earlier, $P(\textbf{Service})>\beta$ since the first honest block contains $tx$ and becomes part of the offset chain as soon as it is mined with probability $\beta$. If one were to consider pre-mining strategies, $P(\textbf{Double-Spending})$ could be increased at the expense of $P(\textbf{Service})$, hence security threats can be even more imminent than we display here.
Next, in Figure 4, we provide the probabilities of the events Double-Spending and Service for each attack cycle of $(k+1)$-stealth mining with $k=6$ and the normalized revenue ratio. Again, at any point above and right of the red curve, a transaction is at the risk of getting replaced by a conflicting transaction after being confirmed at no-cost to an adversary. The associated risk at each attack cycle is shown in Figure 4(a). We note that compared to selfish mining, in stealth mining, for a given $\alpha$, increasing $\gamma$ increases the probability of Double-Spending as we were aiming. However, in stealth mining, for any given $\alpha$ and $\gamma$, $P(\textbf{Service})$ is larger compared to selfish mining as an attack cycle cannot end with Move-Funds event in stealth mining. Essentially, in addition to the revenue decrease compared to the $L$-stubborn mining, this is the other cost of performing the attack by stealth and increasing $P(\textbf{Double-Spending})$.
@@
For $(k+1)$-stubborn mining, the double-spending risk for $k=6$, $\alpha=0.41$ and $\gamma=0$ is $0.092$. For $k=6$ and $\gamma=0$, as $\alpha$ approaches $0.5$, the double-spending risk approaches $0.21$. On the other hand, for $k=6$, $\alpha=0.41$ and $\gamma=1$, the double-spending risk is $0.002$, whereas service probability is $0.59$. For $\gamma=0$, the same values apply to $(k+1)$-stealth mining as well. On the other hand, for $(k+1)$-stealth mining with $k=6$, $\alpha=0.41$ and $\gamma=1$, the double-spending risk is increased to $0.108$, whereas service probability increases to $0.892$. As mentioned earlier, the cost is a cut in revenue ratio, which drops from $\frac{\rho_{7}}{\alpha}=1.639$ to $\frac{\sigma_{7}}{\alpha}=1.09$.
In Figure 5, we provide $R^{*}_{k}$ for each $\alpha$ and $\gamma$ value for Bitcoin, where $k=6$, for both stubborn mining and stealth mining attacks defined earlier. The white region is the region where $R^{*}_{k}=0$, i.e., the revenue ratio is already larger than honest mining. As the modification we provided for the attack resulted in larger $P(\textbf{Double-Spending})$ in stealth mining, $R^{*}_{k}$ value is quite low even for $\alpha$ and $\gamma$ values with $\frac{\sigma_{k+1}}{\alpha}<1$. On the other hand, it seems for $\alpha<0.2$, the double-spending rewards need to be extremely large for both of the attacks to be profitable.
7 Conclusion